OpenStudy (anonymous):

Consider the function f(x) = x2 − 2x + 6. Find the value of f(x) when x = 3 and select the correct answer below. 1 3 6 9

Vocaloid (vocaloid):

x = 3 replace "x" with 3 f(x) = x^2 - 2x + 6 f(3) = (3)^2 - 2(2) + 6 = ?

OpenStudy (anonymous):

11

Vocaloid (vocaloid):

*small correction f(3) = (3)^2 - 2(3) + 6 = ?

OpenStudy (anonymous):

9

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

can you answer another one?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

Which of the following expressions represents a function? {(1, 2), (4, –2), (8, 3), (9, –3)} y2 = 16 − x2 2x2 + y2 = 5 x = 7

Vocaloid (vocaloid):

^ I made a mistake when I was typing that out the second term should be 2(3) not 2(2)

OpenStudy (anonymous):

ok can you solve the question igabe you still?

Vocaloid (vocaloid):

the first one is a function, since it gives a unique y value for each input value

OpenStudy (anonymous):

thanks